On another planet, a marble is released from rest at the top of a high cliff. It falls 4.00 m in the first 1 s of its motion. Through what additional distance does it fall in the next 1 s?
1
Expert's answer
2020-10-05T10:54:52-0400

Solution

Displacement in 1sec can be given by second equation of motion

So

"S_1=ut+\frac{1}{2}at^2"

And

S1=4m t=1sec u=0m/sec

So

"a=\frac{2s}{t^2}\\a=\frac{2\times 4}{1^2\\}\\a=8m/s^2"

So now displacement in 2sec

"S_2=\frac{1}{2}at_2^2"


"S_2=\frac{1}{2}\times8\times2^2\\=16m"

Displacement t=1 and t=2 sec time interval

"S_3=S_2-S_1"

"S_3=16-4=12m"

So additional distance traveled in the next 1sec is 12m .